a swan on a lake gets airborne by flapping its wings and running on top of the water.

etheexpertta.com-tracking id: 3475-8B-7C-45-A432-17049. In accordance with Expert TAs Terms of Service. copying this i sharing website is strictly forbidden. Doing so may result in termination of your Expert TA Account. rt (a) If the swan must reach a speed of 4.5 m/s to take off and it accelerates from rest at a constant rate of 0.35 m/s-, how far, in meters will it travel before becoming airbome? Grade Summary Deductions 0 Potential 1009 tan() | π| ( cotan0 asin0acos0 atan0 acotansinh0 cosh0 tanh0cotanh( Attempts remaining (3% per attempt) detailed view 1 2 3 END Degrees Radians VO DE I: CEE Submit Hint I give up! Hints: 0% deduction per hint. Hints remaining: 2 Feedback: 0% deduction per feedback. ma 50% Part (b) How long will that take, in seconds?
Answer #1

a)

Since it starts from rest ,initial velocity

Vo=0

From

V2=Vo2+2ad

4.52=0+2*0.35*d

d=28.93 m

b)

From

V=Vo+at

4.5=0+(0.35)t

t=12.857 s